Warner company purchased twenty-eight units of a product for $19 each and later purchased fourteen more for 19.50 each. If the company uses the weighted average cost flow method, and it sold one unit of the product for $35, its gross margin would be $15.83. TRUE OR FALSE?
Answer #1

Total cost = (28*19) + (14*19.50) = 805

Total units = 28+14 = 42

Average cost per unit = 805/42 = 19.17 per unit

Gross profit = Sales per unit - Cost per unit

= 35 - 19.17

= 15.83

Hence true